Session: Feb 2017 (session 1)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Robledo (Cristina Garcia Miro at Wed Mar 08 08:15:14 2017) Failure: ANT
After "Follow the Sun" DSN installation and testing antenna controller was declared red. Not possible to observe.

Zelenchukskaya (Lysenkova M. at Wed Mar 08 12:09:54 2017) Success:

Badary (Zhiritsky at Wed Mar 08 12:10:08 2017) Success:

Svetloe (Isaenko O. at Wed Mar 08 12:11:03 2017) Minor failures: RECO
Scans No0001-No0007 missed due to Mark5B+ problems.

Noto (Nocita at Wed Mar 08 13:44:36 2017) Success:

Onsala (Jun Yang at Wed Mar 08 14:05:41 2017) Success:

Effelsberg (Uwe at Wed Mar 08 14:54:07 2017) Minor failures: WIND
Only some 30 minutes lost at the end of the observations because of wind stop at 11:30 UT.

Urumqi (Peng Li at Wed Mar 08 15:23:29 2017) Failure:
Because the moon observation, not observed in this experiment

Yebes (Daniel Valero at Thu Mar 09 07:30:58 2017) Minor failures:
dbbc firmware didn't work properly at bbc1 and bbc3 until scan no0016

Medicina (Roda at Thu Mar 09 08:49:38 2017) Success:

Effelsberg (Uwe at Fri Mar 10 07:52:21 2017) Success:
Some additional notes on the weather: heavy rain and windy most of the time and therefore higher Tsys than usual.

Noto (P.Cassaro at Sat Mar 11 10:47:34 2017) Success:

Onsala (Jun Yang at Tue Mar 14 15:21:28 2017) Success:
Light snow at the beginning of the observations.

Torun (Pawel Wolak at Fri Mar 17 11:10:52 2017) Success: